a young man 150cm tall cast a shadow that was 200cm long. his father's shadow cast at same time of the day was 240cm long how tall was his father?
Found 2 solutions by natroa08, checkley77:
Answer by natroa08(4)

this is how you would set up the proportion
150................X
⎯⎯⎯ = ⎯⎯⎯ (sorry ignore the dots)
200................240
to solve for x you simply just cross multiply like this:
150 x 240= 36000
then you divide 36000 by 200
and that gives you 180 cm. and there's your answer!

Answer by checkley77(12844)

240/200=X/150
200X=150*240
200X=36,000
X=36,000/200
X=180 CM IS THE HEIGHT OF THE FATHER.